A solution is prepared by mixing 100 mL of 2M NaCl, 100 mL of 1M CaCl2 and 650 mL of water. What is the molar ratio of Ca:Cl:Na in the final solution?

please show full solution, (Indicate your answers to two digits after the decimal)
Answer #1

no of moles of NaCl = molarity x volume in liters = 2 x 0.1 = 0.2moles of NaCl

no of moles of CaCl2 = 1x 0.1 = 0.1 moles

one mole of NaCl will have one mole of Na+ and one mole of Cl-

0.2 moles of NaCl will have 0.2 moles of Na+ and 0.2 moles of Cl-

0.1 moles of CaCl2 will have 0.1 moles of Ca2+ and 2 x 0.1 = 0.2 moles of Cl- ions

to tal volume = 100 + 100 + 650 = 850 mL

Molarity of Ca = no of moles of Ca / total volume in liters

= 0.1 / 0.85 = 0.1176

Molarity of Cl = total moles of Cl / total volume

= 0.2 + 0.2 / 0.85 = 0.4 / 0.85 = 0.47 M

Molarity of Na+ = 0.2 / 0.85 = 0.235 M

What is the molar ratio of Ca:Cl:Na = 0.1 : 0.4 : 0.2

= 1: 4: 2